OpenStudy (emetuo89):

the gasoline tank on a van initially read 5/8 full. when 14 gallons were added to the tank, the gauge read 7/8 full. how many more gallons are needed to fill the tank?

OpenStudy (emetuo89):

Could it be 7 more gallons?

OpenStudy (texaschic101):

well...it was 5/8 and after 14 gallons were added, it read 7/8. That means that 14 gallons is 2/8 or 1/4 because 7/8 - 5/8 = 2/8. So if 1/4 = 14 gallons , what does 4/4 (a full tank) equal. Proportion time... (1/4) / 14 = (4/4) / x --- 1/4 to 14 gal = 4/4 to x gal cross multiply (1/4)(x) = (4/4)(14) 1/4x = 14 x = 14 / (1/4) x = 14 * 4 x = 56 check.. 5/8 of 56 = 5/8 * 56 = 35 gallons -- starting gas added 14 making it 49....7/8 of 56 = 7/8 * 56 = 49 gallons leaving 1/8 of the tank left to fill.....1/8 of 56 = 1/8 * 56 = 7 gallons and 49 + 7 = 56 total gallons so yes...the answer is 7 gallons